Operation Summary for March 25, 2011 - March 28, 2011 |
|
There was no Integrated Luminosity for the period from 3/21/11 to 3/28/11 due to the D16-2 magnet replacement. NuMI is off for target replacement. Friday March 25 The day shift began with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE and MTest experiments taking beam, with the Tevatron off for the D16-2 magnet replacement. During the day shift, Booster and Main Injector RF work caused 45 minutes of downtime. During the evening shift, Operations disabled beam to MTest at 8:00 PM. Saturday March 26 The midnight shift began with MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 4:46 AM. During the day shift, Operations delivered beam to MiniBooNE and MTest. During the evening shift, Operations disabled beam to MTest at 6:18 PM. Sunday March 27 The midnight shift began with MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 7:30 AM. During the day shift, Operations delivered beam to MiniBooNE and MTest. During the evening shift, Operations disabled beam to MTest at 6:04 PM. Monday March 28 The midnight shift began with MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 4:25 AM. The Plans for Monday The plans for today are to deliver beam to MiniBooNE and MTest. |
|
More Information Comments and Suggestions |
| last modified 03/28/2011 email Fermilab |
| Security, Privacy, Legal | |